Statement 1 is a function of demand and supply for goods A.

P = 80 – 2Q
P = 20 + 2Q




i. Find an equilibrium price (RM) and quantity (unit).
ii. Now supplier must pay a tax of RM6 per unit. Find new equilibrium price and quantity.
iii. Draw a graph that shows demand and supply curve before and after taxes.
iv. Calculate the amount of tax passed by consumers.
v. Who have do absorb most of the tax burden?
1
Expert's answer
2021-10-10T16:31:20-0400

i

"Qd=Qs\\80-2Q=20+2Q\\80-20=2Q+2Q\\60=4Q\\Q=15\\P=50"


ii

"Qd=Qs\\80-2Q=20+2Q+6\\80-26=2Q+2Q\\54=4Q\\Q=13.5\\P=53"

iii



iv

"(50\u00d715)-(13.5\u00d753)=34.5"


v. Consumer suffer from the tax burden